we have a u-shaped conducting its side with the open part of the u the meter conducting bar is to slide left or right on the track There is maaneoe o.06 out are moving the bar to the right with a constant speed 400 meters per second Calculate the EMF induced in the circuit when the bar is 1.5 meters from the right of the sideways U 18 V 6) A solenoid has a radius o.125 meters, length 1.8 meters, and 60o turns of wire. If the current in the solenoid is 35 amps, calculate the magnitude of the magnetic field the 0.020 Tesla 0.015 Tesla 2.0 E-7 Tesla 6 E-19 Tesla 2.8 E-5 Tesla 7) A solenoid is setup with the x axis as the central axis. The current in the solenoid is A positive when viewed from the left the solenoid (from negative x axis). Assume a toward charge is located in the middle of the solenoid with a velocity directed up of the top of the page (positive y axis). At this moment, determine the direction magnetic force acting on the positive charge. to the left (-x) to the bottom of the page (-y) out of the page (+z) there is no force on the charge to the right (+x) into the page (-z)
Answer #1

Emf = BLv

Where B is magnetic field

L is length

v is velocity

Substituting values

emf = 0.06*0.75*400 = 18V

b) Use B = uoNI/L to get B